# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ V-E Can DPO improve LLM as a Judge?
Wang et al. [131] explore enhancing the auto-evaluation capabilities of LLMs by training them as generative judges using both positive and negative data through preference optimization. The authors introduce a novel framework—SFR-Judges—that leverages three distinct data collection methods, including chain-of-thought critiques, classification-based judgments, and response deduction, to generate comprehensive preference pairs across various evaluation tasks (single rating, pairwise comparison, and classification). By integrating Direct Preference Optimization (DPO) with supervised fine-tuning (SFT) loss, the generative judge model robustly outperforms strong baselines such as GPT-4o and specialized judge models, achieving superior performance on 10 out of 13 benchmarks. Notably, the approach mitigates common biases like position and length bias, adapts flexibly to diverse evaluation protocols, and provides insightful language feedback for improving downstream generative models. Empirical results further demonstrate that models evaluated with SFR-Judges not only reach a 44% performance on the Alpaca Eval but also achieve an average pairwise comparison accuracy of around 81%, underscoring this method’s promise in advancing LLM evaluation strategies.
V-F SFT Memorizes and RL Generalizes
@@
Traditional supervised fine-tuning (SFT) has long been the default approach for specialized model tuning, but its limitations are becoming increasingly apparent. In a groundbreaking study by researchers from DeepMind, UC Berkeley, NYU, and HKU [132], it was demonstrated that SFT often traps models in a cycle of memorization, significantly degrading performance on out-of-domain tasks—by as much as 79% in some cases. This creates a major challenge: how can models be tuned to not only recall training data but also to adapt, reason, and generalize to varied, real-world scenarios, such as those encountered in enterprise environments?
@@
Enter Reinforcement Fine-Tuning (RFT). Unlike SFT, RFT leverages outcome-based reward signals and reinforcement learning techniques (using PPO, for example) to encourage models to learn from their interactions with evaluation environments. In the study [132], a duplicate Llama 3.2 11B base was first fine-tuned with supervised methods and then refined via reinforcement learning. When evaluated on both text-based (GeneralPoints, an arithmetic reasoning card game) and visual navigation (V-IRL) tasks, RFT demonstrated dramatic improvements—boosting performance on out-of-domain tasks by up to 61% and delivering a +34% improvement on the V-IRL task over previous state-of-the-art models.
@@
V-G Does structure prompting impact reasoning?
@@
Tam et al. [133] explores a significant challenge in building real-world LLM applications: balancing structured prompting with the preservation of reasoning abilities. Researchers investigate whether imposing formats such as JSON, XML, or YAML—common in many structured prompts—might undermine an AI’s capacity to reason effectively. The study reveals mixed outcomes by comparing various methods, including constraint decoding, format-restricting instructions, and a two-step natural language-to-format approach across diverse datasets and models. While structured formats can enhance consistency in some cases, they often lead to a decline in reasoning performance. For instance, although specific models perform optimally with JSON, others show notable drops in performance when forced into a rigid schema. This nuanced analysis underscores the complex trade-off between ensuring output uniformity and maintaining the cognitive flexibility required for advanced reasoning tasks.
V-H How effective are LLMs at long-context RAG?
@@
Databricks Mosaic Research [134] systematically investigates the performance of LLMs in long-context Retrieval Augmented Generation (RAG) settings through over 2,000 experiments on 13 open and closed models across four curated datasets. The study reveals that while retrieving additional documents generally enhances RAG performance, most models begin to degrade beyond certain context thresholds—around 32k tokens for models like Llama-3.1-405B and up to 64k tokens for GPT-4. Moreover, different models exhibit distinct failure modes under long-context conditions: for example, Claude 3.5 shows a sharp increase in copyright-related failures from 3.7% at 16k to 49.5% at 64k, and DBRX’s ability to follow instructions deteriorates from 5.2% at 8k to 50.4% at 32k. Some models, such as Mixtral, even generate repeated content, and many suffer from a ”lost in the middle” effect, where information in the central portions of long texts is not effectively utilized. These findings suggest that the optimal context length is highly model- and task-dependent, and that the lack of long-context post-training may be a key factor in the observed performance limitations.
@@
Laban et al. [135] introduces the ”Summary of a Haystack” (SummHay) benchmark, a novel framework designed to evaluate long-context understanding in LLMs and retrieval-augmented generation (RAG) systems. As modern LLMs can handle millions of tokens, traditional tasks like Needle-in-a-Haystack fail to capture the complexity of extended information extraction. In response, the authors synthesize ”haystacks” of documents where key insights are deliberately repeated, challenging models to generate summaries that not only capture these insights (Coverage) but also accurately cite the corresponding source documents (Citation). The evaluation protocol is highly reproducible, automatically scoring summaries on these two critical dimensions. A comprehensive analysis across 10 LLMs and 50 RAG systems reveals that even with Oracle signals for document relevance, current systems lag human performance (56%) by more than 10 points on a joint score. Notably, models such as Google DeepMind Gemini 1.5 pro outperform others—with and without RAG—while long-context LLMs like GPT-4o and Claude 3 Opus score below 20% when not aided by a retriever. Furthermore, smaller models integrated with RAG, such as Claude 3 Haiku or Gemini 1.5 Flash, sometimes outperform larger LLMs, underscoring the importance of effective retrieval strategies in handling long-context tasks. Overall, SummHay provides a challenging yet precise benchmark to drive future improvements in long-context information extraction and enterprise RAG systems.
V-I How important is the reference model in DPO?
@@
Liu et al. [136] investigates a critical yet underexplored aspect of Direct Preference Optimization (DPO) for instruction fine-tuning of large language models, focusing on the role of the reference model or policy. The authors examine three key research questions: the optimal strength of the KL-divergence constraint (controlled by $\beta$), the necessity of a reference model in DPO, and whether a stronger, compatible reference model can enhance performance. Their experiments reveal that reducing the KL-divergence constraint improves DPO performance up to an optimal point (with scores increasing from 12.36 to 16.25), beyond which performance degrades. Moreover, using a stronger reference model—such as one that is already DPO-ed or aligned—further boosts performance (from 16.25 to 20.25), whereas employing an incompatible reference model (e.g., Llama 3) yields suboptimal results. Overall, the study highlights the confounding role of the reference policy in DPO and provides valuable insights for best practices, as demonstrated by an improved Mistral 7B AlpacaEval2 score from 12.36 to 20.25.
V-J Online and offline alignment algorithms
@@
Tang et al. [71] provides an in-depth empirical comparison between online and offline reinforcement learning from human feedback (RLHF) methods for large language model alignment. While offline approaches rely on fixed datasets for preference tuning, online RLHF continuously generates new training data from the evolving policy, leading to a significant performance edge. Using Identity Preference Optimization (IPO) as a common framework, the authors demonstrate that online methods yield higher win rates, largely due to more diverse and effective on-policy data coverage. Intriguingly, the study reveals a trade-off: offline-trained policies excel in pairwise classification but underperform in generative tasks, whereas online-trained models, despite being less adept at classification, generate superior outputs. This performance gap persists across both contrastive and non-contrastive loss functions and remains unmitigated by merely scaling up model size. Overall, the findings underscore the pivotal role of on-policy sampling in RLHF and highlight fundamental challenges in offline alignment algorithms, suggesting that iterative approaches to make offline datasets more on-policy could enhance performance.
V-K Leveraging LLMs for Code Compilation and Optimization
@@
Meta team [137] introduces the Large Language Model Compiler (LLM Compiler), a robust suite of pre-trained models tailored for code and compiler optimization—a domain that has been largely overlooked despite LLMs’ impressive performance in software engineering. Built on the foundation of Code Llama, the LLM Compiler is available in 7B and 13B parameter sizes and has been trained on an extensive corpus of 546 billion tokens of LLVM intermediate representations (IRs) and assembly code from x86_64, ARM, and CUDA. Moreover, through instruction fine-tuning on an additional 164 billion tokens, the models have learned to interpret compiler behavior effectively. Empirical evaluations reveal that the fine-tuned versions can achieve 77% of the optimization potential of an autotuning search and attain a 45% disassembly round trip accuracy with a 14% exact match rate. Released under a bespoke commercial license, this work offers a scalable and cost-effective foundation for both academic research and industrial applications in compiler optimization.
V-L Training MoE architecture
@@
Recent developments in LLMs have increasingly incorporated the Mixture-of-Experts (MoE) paradigm, which extends conventional transformer architectures by replicating feedforward blocks into multiple specialized ”experts.” A dynamic routing mechanism in this framework allocates input tokens to select experts, thereby significantly increasing model capacity (ranging from 20B to 600B parameters) without a proportional increase in per-token computational operations. This architectural modification theoretically enhances model intelligence while preserving latency, albeit at the cost of elevated memory usage. The inherent challenges of MoE, particularly the need for efficient multi-GPU parallelization and optimized inter-device communication, have been addressed by only a limited number of codebases, including DeepSpeed [138], MegatronLM [139], and MosaicML’s LLM Foundry [140]. In this context, the recent open-source release from DeepSeek [141], which includes a state-of-the-art communication and orchestration library with advanced FP8 support, represents a significant contribution to the field. This release underpins the development of DeepSeek-R1—a state-of-the-art MoE model noted for its impressive performance, training efficiency (with an estimated training cost of \$6M), and optimized inference capabilities—and provides valuable insights into efficient MoE parallelism.
VI Conclusion
@@
Recent advances in reasoning LLMs, exemplified by models such as DeepSeek-R1, OpenAI’s o1 & o3, and GPT-4o—have underscored the value of intermediate multistep processing to solve complex problems ranging from advanced math and coding challenges to intricate puzzles. By generating intermediate reasoning steps, these systems provide transparency into their thought processes, leading to more accurate and reliable outcomes than direct-response language models.
@@
To build and refine these reasoning models, researchers have explored various strategies including inference-time scaling, pure reinforcement learning (as seen in DeepSeek-R1-Zero), combined supervised fine-tuning with reinforcement learning, and pure supervised fine-tuning with distillation. Models like Qwen-32B and Llama-based variants have also been adapted using these techniques, demonstrating that emergent reasoning capabilities can be effectively nurtured through a blend of reinforcement learning and fine-tuning. These approaches enhance performance and help mitigate issues such as overthinking and excessive verbosity during inference.
@@
In this paper, we presented a comprehensive review of the advancements in large language models and their training methodologies that have emerged between 2023 and 2025. Our analysis spanned a diverse array of state-of-the-art LLMs. It provided an in-depth comparison of their performance across multiple benchmarks, offering valuable insights into their evolving capabilities. We examined innovative training strategies—including a mixture of experts, retrieval-augmented generation, and reinforcement learning—as well as architectural improvements that have significantly enhanced model performance.

### v2

@@ VII-E Can DPO improve LLM as a Judge?
Wang et al. [wang2024direct] explore enhancing the auto-evaluation capabilities of LLMs by training them as generative judges using both positive and negative data through preference optimization. The authors introduce a novel framework SFR-Judges that leverages three distinct data collection methods, including chain-of-thought critiques, classification-based judgments, and response deduction, to generate comprehensive preference pairs across various evaluation tasks (single rating, pairwise comparison, and classification). By integrating Direct Preference Optimization (DPO) with supervised fine-tuning (SFT) loss, the generative judge model robustly outperforms strong baselines such as GPT-4o and specialized judge models, achieving superior performance on 10 out of 13 benchmarks. Notably, the approach mitigates common biases like position and length bias, adapts flexibly to diverse evaluation protocols, and provides insightful language feedback for improving downstream generative models. Empirical results further demonstrate that models evaluated with SFR-Judges not only reach a 44% performance on the Alpaca Eval but also achieve an average pairwise comparison accuracy of around 81%, underscoring this method’s promise in advancing LLM evaluation strategies.
This line of research opens up important avenues for building more interpretable and unbiased evaluation frameworks. One promising direction involves extending judge models to support multi-turn or multi-agent conversations, enabling richer discourse-level assessments. Another is the incorporation of uncertainty quantification or calibration techniques to improve judge confidence and reliability. Moreover, adapting DPO-based judge models for multilingual or domain-specific evaluation tasks such as scientific reasoning or legal analysis could enhance their practical applicability. Finally, integrating generative judge outputs into closed-loop reinforcement learning setups may enable continuous model self-improvement through informed feedback.
VII-F SFT Memorizes and RL Generalizes
@@
Traditional supervised fine-tuning (SFT) has long been the default approach for specialized model tuning, but its limitations are becoming increasingly apparent. In a groundbreaking study by researchers from DeepMind, UC Berkeley, NYU, and HKU [chu2025sftmemorizesrlgeneralizes], it was demonstrated that SFT often traps models in a cycle of memorization, significantly degrading performance on out-of-domain tasks by as much as 79% in some cases. This creates a major challenge: how can models be tuned to not only recall training data but also to adapt, reason, and generalize to varied, real-world scenarios, such as those encountered in enterprise environments?
@@
Enter Reinforcement Fine-Tuning (RFT). Unlike SFT, RFT leverages outcome-based reward signals and reinforcement learning techniques (using PPO, for example) to encourage models to learn from their interactions with evaluation environments. In the study [chu2025sftmemorizesrlgeneralizes], a duplicate Llama 3.2 11B base was first fine-tuned with supervised methods and then refined via reinforcement learning. When evaluated on both text-based (GeneralPoints, an arithmetic reasoning card game) and visual navigation (V-IRL) tasks, RFT demonstrated dramatic improvements boosting performance on out-of-domain tasks by up to 61% and delivering a +34% improvement on the V-IRL task over previous state-of-the-art models.
@@
Future work should explore curriculum-aware reinforcement fine-tuning that gradually shifts from memorization to generalization. Additionally, incorporating meta-learning principles into RFT could allow models to adapt to unseen domains with limited supervision quickly. Another promising direction involves developing diagnostic benchmarks to explicitly measure the trade-offs between memorization and generalization. Finally, expanding RFT to encompass multimodal environments (e.g., vision-language, embodied agents) will be crucial for evaluating its full potential in real-world reasoning settings.
VII-G Does structure prompting impact reasoning?
@@
Tam et al. [tam2024let] explores a significant challenge in building real-world LLM applications: balancing structured prompting with the preservation of reasoning abilities. Researchers investigate whether imposing formats such as JSON, XML, or YAML common in many structured prompts might undermine an AI’s capacity to reason effectively. The study reveals mixed outcomes by comparing various methods, including constraint decoding, format-restricting instructions, and a two-step natural language-to-format approach across diverse datasets and models. While structured formats can enhance consistency in some cases, they often lead to a decline in reasoning performance. For instance, although specific models perform optimally with JSON, others show notable drops in performance when forced into a rigid schema. This nuanced analysis underscores the complex trade-off between ensuring output uniformity and maintaining the cognitive flexibility required for advanced reasoning tasks.
Looking ahead, future research could explore adaptive prompting mechanisms that dynamically adjust structural constraints based on task complexity or reasoning depth. Another direction involves training format-aware models that can decouple syntactic conformity from semantic reasoning, possibly through multi-task objectives. Moreover, fine-tuning or reinforcement learning strategies that explicitly reward both structural correctness and reasoning quality could help bridge the gap. Developing benchmarks that isolate format-induced reasoning degradation will also be crucial for better diagnostic insights.
VII-H How effective are LLMs at long-context RAG?
@@
Databricks Mosaic Research [leng2024long] systematically investigates the performance of LLMs in long-context Retrieval Augmented Generation (RAG) settings through over 2,000 experiments on 13 open and closed models across four curated datasets. The study reveals that while retrieving additional documents generally enhances RAG performance, most models begin to degrade beyond certain context thresholds around 32k tokens for models like Llama-3.1-405B and up to 64k tokens for GPT-4. Moreover, different models exhibit distinct failure modes under long-context conditions: for example, Claude 3.5 shows a sharp increase in copyright-related failures from 3.7% at 16k to 49.5% at 64k, and DBRX’s ability to follow instructions deteriorates from 5.2% at 8k to 50.4% at 32k. Some models, such as Mixtral, even generate repeated content, and many suffer from a ”lost in the middle” effect, where information in the central portions of long texts is not effectively utilized. These findings suggest that the optimal context length is highly model- and task-dependent, and that the lack of long-context post-training may be a key factor in the observed performance limitations.
@@
Laban et al. [laban2024summary] introduces the ”Summary of a Haystack” (SummHay) benchmark, a novel framework designed to evaluate long-context understanding in LLMs and retrieval-augmented generation (RAG) systems. As modern LLMs can handle millions of tokens, traditional tasks like Needle-in-a-Haystack fail to capture the complexity of extended information extraction. In response, the authors synthesize ”haystacks” of documents where key insights are deliberately repeated, challenging models to generate summaries that not only capture these insights (Coverage) but also accurately cite the corresponding source documents (Citation). The evaluation protocol is highly reproducible, automatically scoring summaries on these two critical dimensions. A comprehensive analysis across 10 LLMs and 50 RAG systems reveals that even with Oracle signals for document relevance, current systems lag human performance (56%) by more than 10 points on a joint score. Notably, models such as Google DeepMind Gemini 1.5 pro outperform others with and without RAG while long-context LLMs like GPT-4o and Claude 3 Opus score below 20% when not aided by a retriever. Furthermore, smaller models integrated with RAG, such as Claude 3 Haiku or Gemini 1.5 Flash, sometimes outperform larger LLMs, underscoring the importance of effective retrieval strategies in handling long-context tasks. Overall, SummHay provides a challenging yet precise benchmark to drive future improvements in long-context information extraction and enterprise RAG systems.
Future research could explore hybrid training approaches incorporating long-context conditioning during pre-training and fine-tuning stages. Developing retrieval-aware attention mechanisms or memory-efficient compression strategies could mitigate middle-content loss and repetition. Additionally, co-training retrievers with decoders using joint objectives may yield more coherent and well-grounded summaries. Finally, benchmarks like SummHay could be extended to include multilingual and multimodal contexts, further testing model generalization under real-world complexity.
VII-I How important is the reference model in DPO?
@@
Liu et al. [liu2024understanding] investigates a critical yet underexplored aspect of Direct Preference Optimization (DPO) for instruction fine-tuning of large language models, focusing on the role of the reference model or policy. The authors examine three key research questions: the optimal strength of the KL-divergence constraint (controlled by $\beta$), the necessity of a reference model in DPO, and whether a stronger, compatible reference model can enhance performance. Their experiments reveal that reducing the KL-divergence constraint improves DPO performance up to an optimal point (with scores increasing from 12.36 to 16.25), beyond which performance degrades. Moreover, using a stronger reference model such as one that is already DPO-ed or aligned further boosts performance (from 16.25 to 20.25), whereas employing an incompatible reference model (e.g., Llama 3) yields suboptimal results. Overall, the study highlights the confounding role of the reference policy in DPO and provides valuable insights for best practices, as demonstrated by an improved Mistral 7B AlpacaEval2 score from 12.36 to 20.25.
Building on these insights, future research could investigate adaptive or dynamic reference policies that evolve during training rather than remaining static. Another promising direction is exploring ensemble or interpolated reference models to stabilize preference gradients and reduce brittleness. Additionally, understanding how reference model compatibility impacts multilingual, multimodal, or tool-augmented tasks could further generalize DPO best practices. Finally, theoretical analysis of DPO convergence under varying reference policy conditions would help ground empirical findings in more rigorous guarantees.
VII-J Online and offline alignment algorithms
@@
Tang et al. [tang2024understanding] provides an in-depth empirical comparison between online and offline reinforcement learning from human feedback (RLHF) methods for large language model alignment. While offline approaches rely on fixed datasets for preference tuning, online RLHF continuously generates new training data from the evolving policy, leading to a significant performance edge. Using Identity Preference Optimization (IPO) as a common framework, the authors demonstrate that online methods yield higher win rates, mainly due to more diverse and effective on-policy data coverage. Intriguingly, the study reveals a trade-off: offline-trained policies excel in pairwise classification but underperform in generative tasks, whereas online-trained models, despite being less adept at classification, generate superior outputs. This performance gap persists across both contrastive and non-contrastive loss functions and remains unmitigated by merely scaling up model size. Overall, the findings underscore the pivotal role of on-policy sampling in RLHF and highlight fundamental challenges in offline alignment algorithms, suggesting that iterative approaches to make offline datasets more on-policy could enhance performance.
These insights suggest several avenues for future exploration. One promising direction involves hybrid alignment frameworks that combine offline pre-training with targeted online refinement cycles to bridge the generative-classification performance gap. Another area of interest is the development of proxy metrics or alignment validators to better guide on-policy sampling strategies. Additionally, advancing offline data augmentation methods, such as imitation distillation of on-policy behavior or synthetic preference mining, could help make static datasets more reflective of real-time interaction patterns. Finally, benchmarking alignment techniques under multi-agent, multilingual, or adversarial feedback settings could further elucidate the generalization limits of current alignment algorithms.
VII-K Leveraging LLMs for Code Compilation and Optimization
@@
Meta team [cummins2024meta] introduces the Large Language Model Compiler (LLM Compiler), a robust suite of pre-trained models tailored for code and compiler optimization a domain that has been largely overlooked despite LLMs’ impressive performance in software engineering. Built on the foundation of Code Llama, the LLM Compiler is available in 7B and 13B parameter sizes and has been trained on an extensive corpus of 546 billion tokens of LLVM intermediate representations (IRs) and assembly code from x86_64, ARM, and CUDA. Moreover, through instruction fine-tuning on an additional 164 billion tokens, the models have learned to interpret compiler behavior effectively. Empirical evaluations reveal that the fine-tuned versions can achieve 77% of the optimization potential of an autotuning search and attain a 45% disassembly round trip accuracy with a 14% exact match rate. Released under a bespoke commercial license, this work offers a scalable and cost-effective foundation for both academic research and industrial applications in compiler optimization.
Future research could investigate integrating LLM-based compilers with traditional optimization pipelines to create hybrid systems that combine statistical reasoning with symbolic guarantees. Another promising direction involves training multilingual or cross-architecture compiler models that generalize across IRs and assembly code from various platforms. Additionally, fine-tuning LLM compilers using reinforcement learning with performance-based rewards could further push the boundaries of learned optimization. Finally, benchmarking compiler-oriented LLMs on real-world codebases and resource-constrained edge platforms would offer deeper insights into their practical viability and robustness.
VII-L Training MoE architecture
@@
Recent developments in LLMs have increasingly incorporated the Mixture-of-Experts (MoE) paradigm, which extends conventional transformer architectures by replicating feedforward blocks into multiple specialized ”experts.” A dynamic routing mechanism in this framework allocates input tokens to select experts, thereby significantly increasing model capacity (ranging from 20B to 600B parameters) without a proportional increase in per-token computational operations. This architectural modification theoretically enhances model intelligence while preserving latency, albeit at the cost of elevated memory usage. The inherent challenges of MoE, particularly the need for efficient multi-GPU parallelization and optimized inter-device communication, have been addressed by only a limited number of codebases, including DeepSpeed [deepspeed-moe], MegatronLM [megatron-lm], and MosaicML’s LLM Foundry [llm-foundry]. In this context, the recent open-source release from DeepSeek [deepep], which includes a state-of-the-art communication and orchestration library with advanced FP8 support, represents a significant contribution to the field. This release underpins the development of DeepSeek-R1 a state-of-the-art MoE model noted for its impressive performance, training efficiency (with an estimated training cost of \$6M), and optimized inference capabilities and provides valuable insights into efficient MoE parallelism.
Future work could explore intelligent expert routing strategies beyond token-level gating to consider task context, domain specialization, or user intent. Another research direction involves hierarchical or sparse MoE configurations to reduce memory and communication bottlenecks further. Co-optimizing model architecture and training pipelines through neural architecture search (NAS) or reinforcement learning may yield more scalable MoE designs. Finally, extending MoE capabilities to multilingual and multimodal inputs could unlock new frontiers in generalized reasoning and efficient LLM deployment.
VIII Conclusion
@@
Recent advances in reasoning LLMs, exemplified by models such as DeepSeek-R1, OpenAI’s o1 & o3, and GPT-4o have underscored the value of intermediate multistep processing to solve complex problems ranging from advanced math and coding challenges to intricate puzzles. By generating intermediate reasoning steps, these systems provide transparency into their thought processes, leading to more accurate and reliable outcomes than direct-response language models.
@@
To build and refine these reasoning models, researchers have explored various strategies, including inference-time scaling, pure reinforcement learning (as seen in DeepSeek-R1-Zero), combined supervised fine-tuning with reinforcement learning, and pure supervised fine-tuning with distillation. Models like Qwen-32B and Llama-based variants have also been adapted using these techniques, demonstrating that emergent reasoning capabilities can be effectively nurtured through a blend of reinforcement learning and fine-tuning. These approaches enhance performance and help mitigate issues such as overthinking and excessive verbosity during inference.
@@
In this paper, we presented a comprehensive review of the advancements in large language models and their training methodologies that have emerged between 2023 and 2025. Our analysis spanned a diverse array of state-of-the-art LLMs. It provided an in-depth comparison of their performance across multiple benchmarks, offering valuable insights into their evolving capabilities. We examined innovative training strategies, including a mixture of experts, retrieval-augmented generation, and reinforcement learning, as well as architectural improvements that have significantly enhanced model performance.
@@
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.